Fred needs 55 ml of 18% sucrose solution to do his science experiment. He needs to mix a 10% sucrose solution with some 35% sucrose solution to get the 18% solution he wants. How much of each should Fred use? (Include units with your numerical answers.)
Answer by addingup(3677)

0.10x+0.35(55-x) = 0.18(55)
0.10x+19.25-0.35x = 9.9
-0.25x = -9.35
x = 37.4
He will need 37.4ml of the 10% solution, and:
55-37.4 = 17.6ml of the 35% solution